# Differentially Private Zeroth-Order Methods for Scalable Large Language Model Fine-tuning

### v4

@@ 4.4. DP-ZO Stagewise Fine-Tuning with pruning
4.4.1. Constant Pruning Mask
Algorithm 3 is a differentially private ZO fine-tuning framework with constant pruning mask calculated at initial. In Algorithm 3, a new loss function $\mathcal{L}(\boldsymbol{\theta})$ is defined, where $\mathbbm{1}$ is the all ones vector and $|\boldsymbol{\theta}^{[l]}|$ is the element-wise absolute value of parameters in the $l$-th layer. Synaptic saliency scores $\nabla\mathcal{L}_{p}\odot\boldsymbol{\theta}$ are esitimated. The scores will be used in Algorithm 4 to update the vector distribution which helps with convergence.
@@ 8 Update vector distribution $\mathbb{V}$ ;
Pruning can be seen as radically setting the direction of the gradient for frozen parameters to be sampled from $N(0,0)$. We further propose Importance Matrix to guide the training of the remaining parameters. Parameters with high scores are sampled from $N(0,x)(x>1)$ with larger weight, tuned with more effort. The experiment result is shown in Table 2. We propose rank-based important matrix whose standard deviation follows the uniform distribution from A to B based on the rank of parameters remained.
@@ 14                   $M[i][i]=A-\frac{(A-B)\cdot rank}{r\cdot\mathrm{len}(Score)}$
4.4.2. Dynamic Pruning Mask
@@ 4       $\beta_{s}=k\cdot\beta_{s-1},\ T_{s}=2\cdot T_{s-1},\ \eta_{s}=\eta_{s-1}/2$;
DP-ZOPO can be divided into two phases. First, we employ data-free pruning to find the important parameters of $\boldsymbol{\theta}$ (the parameters to be fine-tuned). Next, we use a ZO-based fine-tuning method to optimize the model on $\mathbb{V}$. We denote $\nabla_{\mathbb{V}}f_{\beta}(\boldsymbol{\theta})$ as the zeroth-order gradient on the direction sampled from the distribution $\mathbb{V}$.
Theorem 4.5 (Privacy analysis of DP-ZOPO).
@@
Theorem 4.6.
(18) $\phi_{s}(\boldsymbol{\theta}^{\prime S})-\phi_{s}(\boldsymbol{\theta}^{*})\leq\alpha.$
d The total ZO oracle complexity of DP-ZOPO is
(19) $\mathcal{O}\left(\left(\gamma^{2}+dr_{S}C^{2}+dr_{S}\beta_{S}^{2}L^{4}+\frac{dr_{S}C^{2}P^{2}m\log(P/\delta)}{\epsilon^{2}n^{2}}\right)\cdot\frac{1}{\mu\alpha}\right).$
@@
Compared to the total complexity of DP-ZOSO in Theorem 4.4, the total complexity of DP-ZOPO is reduced by a factor of $r_{S}$ theoretically. In the following Section A.2, DP-ZOPO is also proven to obtain better convergence results than DP-ZOSO empirically.
5. Experiment
We conduct comprehensive experiments on both medium-sized masked LM (RoBERTa-large, 350M (Liu et al., 2019)) and large autoregressive LM (OPT-2.7B (Zhang et al., 2022)) in few-shot settings. We also explore both full-parameter tuning and parameter-efficient fine-tuning like prefix-tuning.
5.1. Experimental Setup
@@
We report the metric of accuracy on downstream tasks that run with random seed=42. For RoBERTa-large, we conduct experiments under the hyperparameters detailed in Table 11. We use the hyperparameters in Table 12 for our experiments on OPT-2.7B. For privacy concerns, we do not use early stopping in our experiments. The final model is saved and tested on 1000 test examples. Datasets: For RoBERTa-large, we consider classification datasets: SST-2(Socher et al., 2013), SST-5(Socher et al., 2013), SNLI(Bowman et al., 2015), MNLI(Williams et al., 2017) and TREC(Voorhees and Tice, 2000). We randomly sample 1000 examples for testing and have 512 examples per class for both training and validation. For OPT experiments, we consider the SuperGLUE dataset collection(Wang et al., 2019) including: CB(De Marneffe et al., 2019), BoolQ(Clark et al., 2019) and MultiRC(Khashabi et al., 2018). We also include SST-2(Socher et al., 2013) in our experiments. We randomly sample 1024 examples for training, 500 examples for validation, and 1000 examples for testing.
@@
Models: We conduct experiments on both masked language model (RoBERTa-large, 350M (Liu et al., 2019)) and autoregressive language model (OPT-2.7B(Zhang et al., 2022)) in few-shot settings. We present our main results in Table 2,3. We include a range of ablation studies, including varying the privacy budget, the pruning strategy and the pruning rate. We also explore both full-parameter tuning and parameter-efficient fine-tuning like prefix-tuning.
@@
Privacy Budgets: We consider various privacy levels with $\epsilon=[2,4,8]$ and dynamic $\delta=1/n$ where $n$ is the number of private data for $(\epsilon,\delta)$-DP. We also include $\epsilon=\infty$ baseline that is trained without adding DP noise. In our experiments, we set the clipping threshold to $C=30$. We include the zero-shot does not incur any privacy loss because we evaluate the pre-trained model directly without fine-tuning on private data.
5.2. Main Results
@@
We conduct experiments with RoBERTa-large on sentiment classification, natural language inference, and topic classification. We follow (Malladi et al., 2023) (Gao et al., 2020) to study the few-shot and many-shot settings, sampling $k$ examples per class for $k=512$. DPZero (Zhang et al., 2023) is tested with moment accountant just for fairness. We summarize the results from Table 2.
@@
Table 2. Experiments on RoBERTAa-large. Our method outperforms zero-shot, DPZero and approaches FT with much less memory.
@@
DP-ZOPO works significantly better than zero-shot and other memory-equivalent methods. On all five diverse tasks, our method can optimize the pre-trained model and consistently perform better than zero-shot and prefix-tuning. We also show for several tasks that DP-ZOPO can outperform DPZero up to 5% even with the same privacy analysis (improve 4.4% on TREC, 5.5% on MNLI with $\epsilon=4$). DP-ZOPO outperforms 1.6% on SST-2 with $\epsilon=4$ compared with the results shown in (Tang et al., 2024).
Pruning improves model performance in private settings. We compare DP-ZOPO (with pruning) and DP-ZOSO (without pruning). We show the performance of DP-ZOPO (static pruning mask) among six different pruning rates in Figure 4. Pruning improves performance greatly in all five tasks in all private settings ($\epsilon=2,4,8$). As a detailed result, the best performance of DP-ZOPO is $74.9\%$ on MNLI, outperforming DP-ZOSO by $7.6\%$ with $\epsilon=4$. In different private settings, the optimal pruning rate varies which will be discussed in Section 5.6. With the promising results from RoBERTa-large, we extend our method to the OPT-2.7B (Zhang et al., 2022). We select SuperGLUE tasks (Wang et al., 2019) (including classification and multiple-choice). We randomly sample 1024 examples for training and 1000 test examples for each dataset.
Table 3. Experiments on OPT-2.7B with privacy budget $\epsilon=4,\delta=1/1024$. ICL: in-context learning; FT-prefix for prefix-tuning. DP-ZOSO and DP-ZOPO performs better than zero-shot, ICL, and DPZero with almost the same memory consumption.
@@
DP-ZOPO works significantly better than zero-shot and ICL. On both classification and multiple-choice tasks, DP-ZOPO exhibits strong performance. On a 2.7B-parameter scale, DP-ZOPO outperforms zero-shot and ICL across all tasks with equivalent memory consumption (details in Table 3). DP-ZOPO outperforms DPZero (Zhang et al., 2023) in all five tasks on autoregressive language OPT-2.7B, we improve the accuracy by $3.9\%$ on BoolQ and $9.2\%$ on CB with $\epsilon=4$. We believe that the quality of data-free pruning can greatly influence optimization and we leave how to enhance the performance of pruning as future work.
@@
Dynamic pruning outperforms Static pruning. In dynamic pruning, we select which parameters to update at each step of training and update only those selected using DP-ZOO. We achieve an improvement of $9.1\%$ on CB and 3.8% on BoolQ with OPT-2.7B. Same conclusions are also drawn on the RoBERTa-large in Table 7. Moreover, incremental pruning strategy is superior to other pruning strategies in most scenarios.
5.3. Memory usage
In this section, we profile the memory usage of zero-shot, ICL, FT, FT-prefix, DP-ZOSO and DP-ZOPO. We test OPT-1.3B and OPT-2.7B with Nvidia A100 GPUs (40GB memory) on SST-2 and report the GPU memory consumption in Table 4.
Table 4. GPU memory comsumption with OPT-1.3B and OPT-2.7B (fine-tuned on SST-2). DP-ZOPO takes $r=1\%$ as the constant pruning rate.
GPU memory (MB) OPT-1.3B OPT-2.7B
Zero-shot 2517.73 5066.17
ICL 2517.73 5066.17
DP-ZOSO 2517.73 5066.17
DP-ZOPO 2737.03 (1.08×1.08$\times$1.08 ×) 5514.65 (1.08×1.08$\times$1.08 ×)
FT-prefix 2528.67 5080.55
FT 7545.08 (3×3$\times$3 ×) 20250.93 (4×4$\times$4 ×)
@@ 5.3. Memory usage
DP-ZOSO costs the same GPU memory consumption as Zero-shot and ICL which only use inference for updates. DP-ZOPO uses pruning mask (the same dimensionality as the model) to freeze model parameters during training which costs quite large GPU memory. However, the pruning rate is usually set quite small which means the pruning mask is a sparse vector. We store non-zero ids of the mask instead of the whole vector to reduce memory consumption overhead. Pruning costs a slight amount of (8%) GPU memory while improving the model utility greatly.
@@
As shown in Table 4, DP-ZOPO exhibits almost the same (1.08×1.08$\times$1.08 ×) memory consumption which offers memory savings of up to 4 times compared to standard FT. Moreover, if we pre-prune the network and update the remaining parameters using DP-ZOO, the pruning mask can be calculated offline and stored locally, without incurring additional GPU memory overhead.
5.4. Dynamic ZO scale parameter
@@
We fine-tune RoBERTa-large on SST-2 with both fixed and dynamic ZO scale parameter to demonstrate the superiority of dynamic ZO scale parameter. We present the final training loss of the fine-tuned model with different ZO scale parameter in Table 5.
@@
Table 5. Training loss of RoBERTa-large on SST-2 with both fixed and dynamic ZO scale parameter. Dynamic ZO scale parameter is reduced during stages. We denote M1 as the scale parameter decreasing from $41e-5\sim 1e-$ 4, M2 as $51e-6\sim 1e-$ 5 and M3 as $41e-6\sim 1e-$ 4.
@@
All experiments are done under $\epsilon=4$ with 6k steps. M3 represents the algorithm with ZO scale parameter increasing from $61e-$ 6 to $41e-$ 4 as the stage advances. It is shown in Table 5 that dynamic ZO scale parameter method M3 ($41e-6\sim 1e-$ 4) exhibits the lowest loss on the training set compared with both fixed ZO scale parameter $41e-$ 4 and $61e-$ 6. Dynamic M2 and M3 both perform better than all fixed ZO scale parameter, further elaborating on the superiority of the dynamic ZO scale parameter approach. For the following experiments of RoBERTa-large, we set the dynamic ZO scale parameter from $61e-$ 6 to $41e-$ 4.
5.5. Zeroth-Order Pruning VS First-Order Pruning
In this section, we compare the effectiveness of pruning in both zeroth-order and first-order methods. In first-order method, the gradient is calculated by backpropagation, and only $r$ of the parameters are updated. First-order method in conjunction with pruning reduces the scale of DP noise but also diminishes the model’s utility since only a subset of parameters is fine-tuned.
Figure 3. Results of fine-tuned RoBERTa-large on SNLI with zeroth-order and first-order method. In zeroth-order method, fine-tuning with pruning helps with optimization under all private settings.
In Figure 3, we find that in first-order method, pruning helps with fine-tuning when the privacy budget is small and does not significantly harm fine-tuning, only reducing the accuracy by $0.2\%$ with larger privacy budget. However, in zeroth-order method, pruning not only scales down the DP noise but also diminishes the gradient approximation error as the dimensionality decreases. Under all private settings ($\epsilon=2,4,8,\infty$), pruning helps with zeroth-order fine-tuning and improves 2.7% when $\epsilon=2$ (more detailed numbers in Table 14). The success of the zeroth-order method with pruning motivates us to investigate the impact of different pruning strategies on DP zeroth-order fine-tuning. We will present more experiment results of different pruning strategies in the following section.
5.6. Data-free Pruning with static mask

### v5

@@ IV-D DP-ZO Stagewise Fine-Tuning with pruning
IV-D1 Constant Pruning Mask
Algorithm 3 is a differentially private ZO fine-tuning framework with constant pruning mask calculated at initial. In Algorithm 3, a new loss function $\mathcal{L}(\boldsymbol{\theta})$ is defined, where $\mathbbm{1}$ is the all ones vector and $|\boldsymbol{\theta}^{[l]}|$ is the element-wise absolute value of parameters in the $l$-th layer. Synaptic saliency scores $\nabla\mathcal{L}_{p}\odot\boldsymbol{\theta}$ are estimated. The scores will be used in Algorithm 4 to update the vector distribution which helps with convergence.
@@ 8 Update vector distribution $\mathbb{V}$ ;
Pruning can be seen as radically setting the direction of the gradient for frozen parameters to be sampled from $N(0,0)$. We further propose Importance Matrix to guide the training of the remaining parameters. Parameters with high scores are sampled from $N(0,x)(x>1)$ with larger weight, tuned with more effort. The experiment result is shown in Table II. We propose rank-based important matrix whose standard deviation follows the uniform distribution from A to B based on the rank of parameters remained.
@@ 14                   $M[i][i]=A-\frac{(A-B)\cdot rank}{r\cdot\mathrm{len}(Score)}$
IV-D2 Dynamic Pruning Mask
@@ 4       $\beta_{s}=k\cdot\beta_{s-1},\ T_{s}=2\cdot T_{s-1},\ \eta_{s}=\eta_{s-1}/2$;
DP-ZOPO can be divided into two phases. First, we employ data-free pruning to find the important dimensions of parameters $\boldsymbol{\theta}$ (the parameters to be fine-tuned). Next, we use a ZO-based fine-tuning method to optimize the model on distribution $\mathbb{V}$. We denote $\nabla_{\mathbb{V}}f_{\beta}(\boldsymbol{\theta})$ as the zeroth-order gradient on the direction sampled from the distribution $\mathbb{V}$.
Theorem 3 (Privacy analysis of DP-ZOPO).
@@
Theorem 4.
In DP-ZOPO, suppose assumption 1 holds and loss function $f(\boldsymbol{\theta},x)$ is $\rho$-weakly-convex of $\boldsymbol{\theta}$. The dimensionality of parameters to be updated in stage $s$ is $d\cdot r_{s}$. Then by setting $\eta_{s}$ = $\alpha_{s}$$\cdot\min\{\frac{Pm}{7\gamma^{2}}$, $\frac{Pme}{56dr_{S}C^{2}}$, $\frac{\epsilon^{2}n^{2}}{7dr_{S}c_{2}^{2}C^{2}PT\log(P/\delta)}$, $\frac{Pm}{448dr_{S}\beta_{S}^{2}L^{4}}\}$ and $\lambda=\frac{7}{2\mu}$, $\eta_{s}T_{s}=\frac{7}{2\mu}$, after $S=\lceil\log(\alpha_{0}/\alpha)\rceil$ stages, we have
$\phi_{s}(\boldsymbol{\theta}^{\prime S})-\phi_{s}(\boldsymbol{\theta}^{*})\leq\alpha.$ (18)
The total ZO oracle complexity of DP-ZOPO is
$\small\mathcal{O}\left(\left(\gamma^{2}+dr_{S}C^{2}+dr_{S}\beta_{S}^{2}L^{4}+\frac{dr_{S}C^{2}P^{2}m\log(P/\delta)}{\epsilon^{2}n^{2}}\right)\cdot\frac{1}{\mu\alpha}\right).$ (19)
@@
Compared to the total complexity of DP-ZOSO in Theorem 2, the total complexity of DP-ZOPO is reduced by a factor of $r_{S}$ theoretically. In the following Section A-B, DP-ZOPO is also proven to obtain better convergence results than DP-ZOSO empirically.
V Experiment
We conduct comprehensive experiments on both medium-sized masked LM (RoBERTa-large, 350M [42]) and large autoregressive LM (OPT-2.7b [4]) in few-shot settings. We present the experimental results of dynamic ZO scale parameter in Section V-D, the effectiveness of pruning with zeroth-order and first-order method in Section V-E. In Section V-F and Section V-G, we compare the performance of different pruning strategies. In Section V-H, we report the experiment results on more diverse tasks.
V-A Experimental Setup
We report the metric on downstream tasks that run with random seed=42. For RoBERTa-large, we conduct experiments under the hyperparameters detailed in Appendix A-B Table X. We use the hyperparameters in Appendix A-B Table XI and Table XII for our experiments on OPT-1.3b and OPT-2.7b. We consider diverse tasks like classification, question-answering, translation and summarization.
@@ V-A Experimental Setup
Classification task: We consider datasets: SST-2 [43], SST-5 [43], SNLI [44], MNLI [45] and TREC [46] on RoBERTa-large. We randomly sample 1000 examples for testing and have 512 examples per class for both training and validation. For OPT, we consider the SuperGLUE dataset collection [47] including: CB [48], BoolQ [49] and MultiRC [50]. We also include SST-2 [43] in our experiments. We randomly sample 1024 examples for training, 500 examples for validation, and 1000 examples for testing.
Question-answering task: We consider SQuAD [51], a popular benchmark dataset particularly for machine reading comprehension and question answering systems which consists of questions posed by crowdworkers on a set of Wikipedia articles. We randomly sample 1024 examples for training, 500 examples for validation, and 1000 examples for testing.
Translation task: We focus on the translation direction from French to English and fine-tune with the commonly used Europarl [52] for sentence-level translation. We fine-tune LLMs using examples that include specifically formatted prompts (French: [fr sent] English:) and their corresponding responses ([en sent]). Due to the complexity of text-generation tasks, we use 10000 examples for training and 1000 examples for testing. We use BLEU and COMET [53] as evaluation metrics to access the performance of fine-tuned models.
Summarization task: We consider CNN/Daily Mail corpus [54] and XSum [55]. The CNN/DM dataset primarily focuses on extractive summarization with multi-sentence summaries, while the XSum dataset emphasizes highly abstractive single-sentence summaries. We use the F1 scores of ROUGE-1/2/L/SUM [56] as our evaluation metrics. We randomly sample 1024 examples for training, 500 examples for validation, and 1000 examples for testing.
@@
Models: We conduct experiments on both masked language model (RoBERTa-large, 350M [42]) and autoregressive language model (OPT-1.3b, OPT-2.7b [4]) in few-shot settings. We include a range of ablation studies, including varying the privacy budget, the pruning strategy and the pruning rate. We also explore both full-parameter tuning and parameter-efficient fine-tuning like prefix-tuning.
@@
Privacy Budgets: We consider various privacy levels with $\epsilon=[2,4,8]$ and dynamic $\delta=1/n$ where $n$ is the number of private data for $(\epsilon,\delta)$-DP. We also include $\epsilon=\infty$ baseline that is trained without DP noise. In our experiments, we set the clipping threshold to $C=30$. We include the zero-shot does not incur any privacy loss because we evaluate the pre-trained model directly without fine-tuning on private data.
V-B Main Results
@@
We conduct experiments with RoBERTa-large on sentiment classification, natural language inference, and topic classification. We follow [20] [57] to study the few-shot and many-shot settings, sampling $k$ examples per class for $k=512$. DPZero [1] is tested with moment accountant just for fairness. We summarize the results from Table II.
@@
TABLE II: Experiments on RoBERTAa-large. Our method outperforms zero-shot, DPZero and approaches FT with much less memory.
@@
DP-ZOPO works significantly better than zero-shot and other memory-equivalent methods. On all five tasks, our method optimizes the pre-trained model and consistently performs better than zero-shot and prefix-tuning. We also show for several tasks that DP-ZOPO can outperform DPZero up to 5% even with same privacy analysis (improve 4.4% on TREC, 5.5% on MNLI with $\epsilon=4$). DP-ZOPO outperforms 1.6% on SST-2 with $\epsilon=4$ compared with the results shown in [26].
Pruning improves model performance in private settings. We compare DP-ZOPO (with pruning) and DP-ZOSO (without pruning). We show the performance of DP-ZOPO (static pruning mask) among six different pruning rates in Figure 6. Pruning improves performance greatly in all five tasks in all private settings ($\epsilon=2,4,8$). As a detailed result, the best performance of DP-ZOPO is $74.9\%$ on MNLI, outperforming DP-ZOSO by $7.6\%$ with $\epsilon=4$. In different private settings, the optimal pruning rate varies which will be discussed in Section V-F. With the promising results from RoBERTa-large, we extend our method to the OPT-2.7b [4]. We select SuperGLUE tasks [47] (including classification and multiple-choice). We randomly sample 1024 examples for training and 1000 test examples for each dataset.
TABLE III: Experiments on OPT-2.7b with privacy budget $\epsilon=4,\delta=1/1024$. ICL for in-context learning; FT-prefix for prefix-tuning. DP-ZOSO and DP-ZOPO performs better than zero-shot, ICL, and DPZero with almost the same memory consumption.
@@
DP-ZOPO works significantly better than zero-shot and ICL. On both classification and multiple-choice tasks, DP-ZOPO exhibits strong performance. On a 2.7b-parameter scale, DP-ZOPO outperforms zero-shot and ICL across all tasks with equivalent memory consumption (details in Table III). DP-ZOPO outperforms DPZero [1] in all five tasks on autoregressive language OPT-2.7b, we improve the accuracy by $3.9\%$ on BoolQ and $9.2\%$ on CB with $\epsilon=4$. We believe that the quality of data-free pruning can greatly influence optimization and we leave how to enhance the performance of pruning as future work.
@@
Dynamic pruning outperforms Static pruning. In dynamic pruning, we select which parameters to update at each stage and update those selected using DP-ZOO. We achieve an improvement of $9.1\%$ on CB and 3.8% on BoolQ with OPT-2.7b. Same conclusions are also drawn on the RoBERTa-large in Table XVII. Moreover, incremental pruning strategy is superior to other pruning strategies in most scenarios.
V-C Memory Usage
@@
In this section, we profile the memory usage of zero-shot, ICL, FT, FT-prefix, DP-ZOSO and DP-ZOPO. We test OPT-1.3b and OPT-2.7b with Nvidia A100 GPUs (40GB memory) on SST-2 and report the GPU memory consumption in Figure 4. DP-ZOSO costs the same GPU memory consumption as Zero-shot and ICL which only use inference for updates. DP-ZOPO uses pruning mask (the same dimensionality as the model) to freeze model parameters during training which costs quite large GPU memory. However, the pruning rate is usually set quite small which means the pruning mask is a sparse vector. We store non-zero ids of the mask instead of the whole vector to reduce memory consumption overhead. Pruning costs a slight amount of (8%) GPU memory while improving the model utility greatly.
Figure 4: The GPU memory consumption and running time with OPT-1.3b and OPT-2.7b on SST-2. DP-ZOSO and DP-ZOPO cost less GPU memory consumption and running time.
@@
As shown in Figure 4, DP-ZOPO exhibits almost the same (1.08×1.08$\times$1.08 ×) memory consumption which offers memory savings of up to 4 times compared to FT. Moreover, if we pre-prune the network and update the remaining parameters using DP-ZOO, the pruning mask can be calculated offline and stored locally, without incurring additional GPU memory overhead. DP-ZOPO also runs faster than first-order methods, iterating 6000 rounds with $58.8\%$ time of FT-prefix and $28.5\%$ of FT.
V-D Dynamic ZO scale parameter
@@
We fine-tune RoBERTa-large on SST-2 with both fixed and dynamic ZO scale parameter to demonstrate the superiority of dynamic ZO scale parameter. We present the final training loss of the fine-tuned model with different ZO scale parameter in Table IV.
@@
TABLE IV: Training loss of RoBERTa-large on SST-2 with both fixed and dynamic ZO scale parameter. Dynamic ZO scale parameter is reduced during stages. We denote M1 as the scale parameter decreasing from $41e-5\sim 1e-$ 4, M2 as $51e-6\sim 1e-$ 5 and M3 as $41e-6\sim 1e-$ 4.
@@
All experiments are done under $\epsilon=4$ with 6k steps. M3 represents the algorithm with ZO scale parameter increasing from $61e-$ 6 to $41e-$ 4 as the stage advances. It is shown in Table IV that dynamic ZO scale parameter method M3 ($41e-6\sim 1e-$ 4) exhibits the lowest loss on the training set compared with both fixed ZO scale parameter $41e-$ 4 and $61e-$ 6. Dynamic M2 and M3 both perform better than all fixed ZO scale parameter, further elaborating on the superiority of the dynamic ZO scale parameter approach. For the following experiments of RoBERTa-large, we set the dynamic ZO scale parameter from $61e-$ 6 to $41e-$ 4.
V-E Zeroth-Order Pruning VS First-Order Pruning
In this section, we compare the effectiveness of pruning in both zeroth-order and first-order methods. In FO method, the gradient is calculated by backpropagation, and only $r$ of the parameters are updated. First-order method in conjunction with pruning reduces the scale of DP noise but also diminishes the model’s utility since only a subset of parameters is fine-tuned.
Figure 5: Results of fine-tuned RoBERTa-large on SNLI with zeroth-order and first-order method. In zeroth-order method, fine-tuning with pruning helps with optimization under all private settings.
In Figure 5, we find that in first-order method, pruning helps with fine-tuning when the privacy budget is small and does not significantly harm fine-tuning, only reducing the accuracy by $0.2\%$ with larger privacy budget. However, in zeroth-order method, pruning not only scales down the DP noise but also diminishes the gradient approximation error as the dimensionality decreases. Under all private settings ($\epsilon=2,4,8,\infty$), pruning helps with zeroth-order fine-tuning and improves 2.7% when $\epsilon=2$ (more detailed numbers in Table XVI). The success of the zeroth-order method with pruning motivates us to investigate the impact of different pruning strategies on DP zeroth-order fine-tuning. We will present more experiment results of different pruning strategies in the following section.
V-F Data-free Pruning with static mask